OpenStudy (anonymous):

What is the value of 1a=2b=10c=20d=100e ?

OpenStudy (anonymous):

go tothis : http://telia.hubpages.com/hub/teliamathshelp-

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

a b c d e are specific constant.

OpenStudy (anonymous):

Answer:- un = 2n + 1

OpenStudy (anonymous):

\[U _{n}=2_{n}+1\]

OpenStudy (anonymous):

?? it is not sequnces or series i think.

OpenStudy (anonymous):

It is a ratio problem

OpenStudy (anonymous):

??

OpenStudy (anonymous):

in terms of a, e=a/100 d=a/20 c=a/10 b=a/2

OpenStudy (anonymous):

so they are all inverse ratios of each other. set e = 1. the rest follows.

OpenStudy (anonymous):

depending on what one of those constants are, this is how they're all related to each other...

OpenStudy (anonymous):

Double, multiply by 5, next letter 200f

OpenStudy (anonymous):

Actually, don't know why I answered this, the question doesn't even make sense....